24.
a = 1 + b,
b = 5 − 2a

Thanks for any help.

Answer by ankor@dixie-net.com(22740)

a = 1 + b,
b = 5 − 2a
Using the 1st equation, we can replace a with (1+b) in the 2nd equation
b = 5 - 2(1+b)
b = 5 - 2 - 2b
b + 2b = 5 - 2
3b = 3
b = 1
:
Find a using the 1st equation, replace b with
a = 1 + 1
a = 2
:
:
Check these solutions in the 2nd equation b = 5 - 2a
1 = 5 - 2(2)
1 = 5 - 4